OpenStudy (anonymous):

I need help!!!!! through:(-2,-3), parallel to y=x+4

OpenStudy (anonymous):

any equation parallel to y=x+4 is y=x+c because it passes through (-2,-3) -3=-2+c c=-3+2=-1 required eq. is y=x-1

OpenStudy (anonymous):

Hint:any equation parallel to y=x+4 has same gradient...

OpenStudy (mathmale):

Or, in other words, ... has the same slope. What is the slope of y=x+4?